OpenStudy (anonymous):

What is a linear equation in x whose only solution is 5

OpenStudy (anonymous):

x - 5 = 0

OpenStudy (anonymous):

if you want a function, f(x) = x - 5

OpenStudy (anonymous):

That solution is 0 though? Would 10 - x = 5 work?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Okay, thank you so much ^_^

OpenStudy (anonymous):

10 - x = 5 10 - 5 - x = 0 5 - x = 0 x - 5 = 0, same thing